# Finite Population Identification and

### v2

@@ Abstract
We develop a new approach for quantifying uncertainty in finite populations, by using design distributions to calibrate sensitivity parameters in finite population identified sets. This yields uncertainty intervals that can be interpreted as identified sets, Bayesian credible sets, or frequentist design-based confidence sets. We focus on quantifying uncertainty about the average treatment effect (ATE) due to missing potential outcomes in a randomized experiment, where our approach (1) yields design-based confidence intervals for ATE which allow for heterogeneous treatment effects but do not rely on asymptotics, (2) provides a new motivation for examining covariate balance, and (3) gives a new formal analysis of the role of randomized treatment assignment. We illustrate our approach in three empirical applications.
@@ 1 Introduction
Since the beginning of statistics, there have been two ways to think about populations:
  1. 1.
    They are finite, as in Neyman (1923), or Bowley (1926) who considered “a sample containing $n$ persons or things [that] is selected from a universe containing $N$….” (page 3).
  2. 2.
    Or they are infinite, as in Fisher (1930), or Student (1908) who assumed “a random sample has been obtained from an indefinitely large* population…. *Note that the indefinitely large population need not actually exist” (page 302).
Inference in finite populations came to be called “design-based inference” or “randomization inference” while inference in infinite populations came to be called “model-based inference” or “super-population inference” (e.g., Wu and Thompson 2020). Econometric theory has traditionally focused on super-population inference, going back to Haavelmo (1944) who explained that “when we describe $s$ as a random variable with a certain probability distribution for each fixed set of values of the variables $x$, we are thinking of a class of hypothetical, infinite populations…” (page 51). Likewise, much of statistics has also focused on the super-population approach, with survey sampling theory as the main exception (see Little 2004). [^3] These distinctions and jargon come from the survey sampling literature (e.g., Särndal et al. 1978), which is how we use them here. In economics, some researchers now use the term “design-based” in a related but not equivalent sense, to refer to specific identification strategies (e.g. Card 2022). We focus on randomized experiments in this paper and hence both meanings of “design-based” apply.
However, in the past decade there has been a resurgence of interest in design-based inference in explicitly finite populations (see section 1.2). Much of this work follows a template largely laid out by Neyman (1923). This template has three steps:
  1. 1.
    Describe the finite population.
  2. 2.
    Describe the design, which explains how the data is obtained.
  3. 3.
    Derive properties of the induced design distribution of statistics that depend on the population and the data.
For example, in the context of survey sampling, the population is a set of numbers $\mathbf{Y}=(Y_{1},\ldots,Y_{N})$ (step 1). A classical design is simple random sampling: Sample exactly $n$ units from $\{1,\ldots,N\}$ uniformly at random (step 2). Let $\mathbf{S}=(S_{1},\ldots,S_{N})$ be binary random variables indicating which units were sampled. A classical theorem is that the sample mean (a function of $\mathbf{S}$ and $\mathbf{Y}$) is design-unbiased for the finite population mean when $\mathbf{S}$ is chosen uniformly at random: $\mathbb{E}_{\text{design}}\left[\frac{1}{n}\sum_{i=1}^{N}S_{i}Y_{i}\right]=\frac{1}{N}\sum_{i=1}^{N}Y_{i}$ (step 3). Viewed from the perspective of econometrics, there is a noticeable omission in this template: What about identification? Identification analysis is a key step throughout much of super-population based econometrics (e.g., Lewbel 2019 or Molinari 2020); is any of that analysis still useful in finite populations? Yes, as we argue in this paper. We use ideas from identification theory to construct a new approach for quantifying uncertainty in finite populations, which we call design-based sensitivity analysis. This approach is non-asymptotic, which allows it to be reliably applied in small populations. It yields uncertainty intervals that can be interpreted as (1) identified sets, (2) Bayesian credible sets, or (3) frequentist design-based confidence sets. And it allows researchers to flexibly impose additional assumptions on unobserved values, transparently quantifying the impact of these assumptions on conclusions.
@@
We develop this approach in the context of randomized experiments with a binary treatment, where the goal is to quantify the uncertainty in the average treatment effect (ATE) due to missing potential outcomes (we also briefly discuss sampling uncertainty). In this context, our approach yields three further contributions: (1) Our uncertainty intervals allow for heterogeneous treatment effects, implying that our approach yields design-based confidence intervals for ATE without relying on asymptotics or imposing homogeneous treatment effects. These intervals are not constructed via hypothesis test inversion, but rather are based on ideas from partial identification. (2) We give a new motivation for examining covariate balance. Specifically, by making an assumption that explicitly links covariates to potential outcomes, we show that observed imbalances in covariates have identifying power for the unobserved realized imbalance in potential outcomes, which allows researchers to obtain tighter bounds on ATE. And (3) we give a new formal analysis of the role of randomized treatment assignment.
@@ 1.1 Our approach to quantifying uncertainty in experiments: A practitioner’s guide
Here we briefly sketch our approach, with an eye toward practical take-aways. Consider the population of six units in table 1 (left). Here $Y_{i}(1)$ and $Y_{i}(0)$ denote potential outcomes and $X_{i}\in\{0,1\}$ denotes realized treatment. The true ATE is 0.55, the difference between the average of $Y_{i}(1)$ over $i=1,\ldots,6$ and the average of $Y_{i}(0)$ over $i=1,\ldots,6$. The data, shown in the right table, only reveals half of the potential outcomes, however. Suppose we know that all outcomes must be between 0 and 1. Then without further assumptions, all we can say for sure about ATE is that it is in the set
@@
Finally, in step 3, we recommend combining these two plots as follows: For any desired ex ante probability $1-\alpha\in(0,1)$, use the bottom plot to find the smallest magnitude of imbalance that occurs with at least $100(1-\alpha)$ex ante probability; denote this value by $K(\alpha)$. For example, if we set $0.91-\alpha=$ 0.9 then the figure shows us that $K=18.9$ satisfies $\underline{p}(K)=1-\alpha$ (see the dotted lines). We can then move to the top plot to read off bounds on ATE for this value of the sensitivity parameter, as shown in the dotted lines. In this case the bounds are about $[-6,22]$. Figure 1(b) shows these bounds for all values of $1-\alpha$. We show that these uncertainty intervals have three interpretations:
@@
    Third, for any fixed $1-\alpha$, we show that these bounds are design-based confidence intervals. That is, across repeated re-assignments of treatment, these bounds will contain the true ATE at least $100(1-\alpha)$of the time.
@@ 1.2 Related literature
Our paper relates to a variety of literatures, but for brevity we only discuss the most closely related work, rather than attempt a comprehensive survey. Manski and Pepper (2018) derive identified sets in an explicitly finite population setting. Like them, we also consider the identifying power of bounded variation type assumptions, which restrict just how far potential outcomes can be from each other. The main difference is that they focus on a setting with observational data whereas we consider randomized experiments. Our focus on experiments allows us to use randomization itself to calibrate the magnitude of the bounded-variation sensitivity parameter, which is the design-based sensitivity analysis we develop in this paper. Like us, Greenland and Robins (1986) also describe the problem of drawing conclusions in finite populations as an identification problem which must be solved by making assumptions about unobserved values of variables. They formally show how exact balance assumptions yield point identification. They then informally explain that “if we randomize…when both samples are large, random differences will in probability be small” (page 415). Many other papers make similar informal remarks, like Lindley (1980, page 590) and Royall and Pfeffermann (1982, page 480). Relatedly, the distinction between ex ante and ex post balance is well known; Greenland (1990) gives a particularly clear discussion, where he considers a randomized experiment with $N=2$. Greenland and Robins (2009) more recently discuss this distinction, and conclude that “randomization (or more generally, ignorability) does not impose “no [ex post] confounding”…rather, it provides…a randomization-based (“objective”) derivation of a prior…that applies after allocation as well as before, and becomes more narrowly centered around zero as the sample size increases. This is a key post-allocation benefit of randomization” (page 5, emphasis in the original). That paper as well as the rest of this prior literature, however, is largely verbal discussion. One of our main contributions is to use those informal ideas to provide a new, detailed, and formal approach to quantifying uncertainty in small finite populations.
@@
Finally, many recent papers on causal inference study explicitly finite populations, including Li and Ding (2017), Ding et al. (2017), Aronow and Samii (2017), Ding (2017), Athey et al. (2018), Kang et al. (2018), Abadie et al. (2020, 2023), Hong et al. (2020), Rambachan and Roth (2020), Wu and Ding (2021), Eckles et al. (2020), Imbens and Menzel (2021), Zhao and Ding (2021), Sävje (2021), Bojinov et al. (2021), Xu (2021), Xu and Wooldridge (2022), Athey and Imbens (2022), Roth and Sant’Anna (2023), Pollmann (2023), Wooldridge (2023), Startz and Steigerwald (2023), Borusyak et al. (2024), and Sancibrián (2024), among many others. Also see Imbens and Rubin (2015) and Ding (2024) for book level surveys. This literature largely follows the template of Neyman (1923) that was discussed earlier, and does not do explicit identification analysis.
@@ 2 Finite Population Identification
We first define the finite population identified set. We then use this definition to derive identified sets for finite population average treatment effects under various assumptions.
@@ 2.1 The finite population identified set
Consider the standard potential outcomes model with a binary treatment. There are four components required to define an identified set:
1. The population: Suppose there are $N$ units in our population. Let =$\mathcal{I}=\{1,\ldots,N\}$ be the set of indices for these units. Each unit $i\in\mathcal{I}$ is associated with the vector of numbers $(Y_{i}(1),Y_{i}(0),X_{i},W_{i})$, where $Y_{i}(1)$ and $Y_{i}(0)$ are potential outcomes, $X_{i}\in\{0,1\}$ is a realized binary treatment, and $W_{i}$ is a $d_{W}$-vector of covariates. The population is the $N\times(3+d_{W})$ matrix of all of these numbers. Let $\mathbf{P}^{\text{true}}\coloneqq(\mathbf{Y}(1),\mathbf{Y}(0),\mathbf{X},\mathbf{W})$ denote this matrix. We use $\mathbf{P}$ to denote alternative possible values of this population matrix. Table 1 shows an example population.
2. Assumptions: We do not observe $\mathbf{P}^{\text{true}}$. We will make assumptions about it, however. Formalize these assumptions as the restriction that $\mathbf{P}^{\text{true}}\in\mathcal{P}$ where $\mathcal{P}$ is a known set of $N\times(3+d_{W})$ matrices. We give examples of $\mathcal{P}$ in section 2.2.
@@
4. The data: Finally, we must describe the data that is observed to the econometrician. To allow for sampling, let $\mathbf{S}=(S_{1},\ldots,S_{N})$ be a vector of indicators where $S_{i}=1$ if unit $i$ appears in the dataset, and zero otherwise, as in section 1. Let $\mathbf{Y}=(Y_{1},\ldots,Y_{N})$ where $Y_{i}\coloneqq Y_{i}(1)X_{i}+Y_{i}(0)(1-X_{i})$ for all $i\in\mathcal{I}$. Then $\mathbf{P}^{\text{data}}$ is the matrix of all rows of $(\mathbf{Y},\mathbf{X},\mathbf{W})$ with $S_{i}=1$. Formalize this construction via the function $\text{MakeData}(\mathbf{P},\mathbf{S})$, so that $\mathbf{P}^{\text{data}}=\text{MakeData}(\mathbf{P}^{\text{true}},\mathbf{S})$. [^4] We define the population at the time period after which treatment has been assigned but before units are sampled. This is standard in the identification literature (e.g., Manski 2003, ch. 7). One could modify the notation to further distinguish population values of potential outcomes and covariates from realized treatment assignment, as none of our concepts or procedures depend on this particular notational choice, but at the cost of additional notation throughout, especially since we mainly focus on the case where all units are sampled.
@@
This definition follows the usual description of the identified set as “the set of parameters…that are consistent with the model and the data” (Tamer 2010, page 184), but with two variations to clarify what is meant by “data” when the population is finite: First, even when all units are sampled, the data is defined by a finite dimensional matrix rather than a joint distribution of random variables (as in def. 3.1 on page 5324 of Matzkin 2007, for example; also see our appendix LABEL:sec:popDistributions for a related discussion). Second, our definition allows for data on some units to be missing (i.e., not sampled), by setting $S_{i}=0$. This is perhaps unusual, since uncertainty due to sampling is traditionally separated from other types of uncertainty, as in Koopmans’ (1949, page 132) original definition of identification. Nonetheless, as we show in this paper, definition 1 provides a useful foundation to quantify uncertainty from both missing units and missing potential outcomes. In particular, $\Theta_{I}$ has two key features: (1) It can always be computed with the data the analyst actually has at hand, and (2) like traditional definitions of identification, it is guaranteed to contain the true parameter of interest, so long as the model is not false. We will revisit these properties below.
Finally, note that some authors define the “data” as the design distribution of the observed data, rather than the dataset that is actually observed. This leads to a different definition of identification that can be applied to finite populations; for example, see the definition of “sampling identification” in Florens and Simoni (2021, def 2.1 on page 5). This design distribution typically depends on the entire population matrix and hence is never actually knowable, even if all units are sampled (because not all potential outcomes are observed). Moreover, under this alternative definition of identification, the existence of a design-unbiased estimator for a parameter implies that the parameter is point identified. In contrast, under definition 1, any parameter that depends on at least one unknown value, like a missing potential outcomes, will typically be partially identified, even if there is a design-unbiased estimator for it. In this paper we argue that this feature of our definition leads to a useful approach for quantifying uncertainty in finite populations.

### v3

@@ Abstract
We develop a new approach for quantifying uncertainty in finite populations, by using design distributions to calibrate sensitivity parameters in finite population identified sets. This yields uncertainty intervals that can be interpreted as identified sets, robust Bayesian credible sets, or uniform frequentist design-based confidence sets. We focus on quantifying uncertainty about the average treatment effect, where our approach (1) yields design-based confidence intervals which allow for heterogeneous treatment effects without using asymptotics, (2) provides a new motivation for examining covariate balance, and (3) gives a new formal analysis of the role of randomization. We illustrate our approach in three empirical applications.
@@ 1 Introduction
Many empirical papers use variation across a small number of units to learn about causal effects. For example, in many settings the variation is across units like states or industries. How does identification of causal effects work in such settings, and how should empirical researchers quantify uncertainty? Our paper gives new answers to these questions, focusing on the context of randomized experiments with a finite number of units.
Traditionally, the $N$ units are assumed to be sampled from a hypothetical infinite “super” population, but recently there has been a resurgence in the “design-based inference” literature (reviewed in section 1.2) which studies explicitly finite populations. [^3] The “design-based” versus “super-population” distinction and jargon come from the survey sampling literature (e.g., Särndal et al. 1978), which is how we use them here. In economics, some researchers now use the term “design-based” in a related but not equivalent sense, to refer to specific identification strategies (e.g. Card 2022). We focus on randomized experiments in this paper and hence both meanings of “design-based” apply. With very few exceptions (e.g., Manski and Pepper 2018, Borusyak et al. 2025, and Rambachan and Roth 2025), most papers that study finite populations do not discuss identification and instead focus solely on proving frequentist properties like consistency or confidence interval validity. In contrast, there is a vast and rich literature on identification in infinite populations [^4] See Christensen and Connault (2023), Obradović (2024), Spini (2024), Deaner (2025), Słoczyński (2025), Blandhol et al. (2025), and Caetano et al. (2026) for a few recent examples, among many others; for comprehensive surveys see Matzkin (2007), Manski (2009), Bontemps and Magnac (2017), Lewbel (2019), Chesher and Rosen (2020), Molinari (2020), and Kline and Tamer (2023).; no corresponding literature on identification in finite populations exists. Our first main contribution, therefore, is to develop a formal theory of identification in finite populations.
The literature on identification in finite populations is so sparse that we are not aware of any source which formally defines identification in an explicitly finite population setting (though some references come close). Nonetheless, it is quite common and “conventional” in the prior literature (Heckman and Vytlacil 2007, footnote 6, page 4786) to informally state that a parameter is point identified if there is a known mapping from the distribution of the data induced by randomization (of treatment assignment, for example) into the parameter. We state this prior definition formally, and then argue that it has two major problems.
First, we show that this prior conventional definition relies on hypothetical data that is impossible to observe, since it involves averaging over counterfactual worlds that are incompatible with each other. As a consequence, unlike in infinite populations, we prove that this prior conventional definition in finite populations implies the values of every potential outcome for every unit are point identified in any setting where the probability of each treatment value is positive (a type of overlap restriction), including non-experimental settings where treatment assignment depends almost arbitrarily on potential outcomes (this finding builds on ideas in Imbens 2018 and Chen et al. 2026, as we discuss in section 2.4). For example, according to the prior conventional definition, assumptions like unconfoundedness or instrument exogeneity are unnecessary to obtain point identification of treatment effects, so long as an overlap restriction holds. Therefore, we argue that the prior conventional definition is not adequate for assessing how informative the actual data one has are about treatment effects.
Second, the prior conventional definition of identification, as well as the broader literature on finite population design-based inference, pre-supposes that there is some kind of known a priori randomization. For example, that randomization may involve treatment assignment, an instrument assignment, or sampling, depending on the setting. These may be plausible in randomized experiments or designed surveys, but in many observational settings it is not clear that the variable of interest (like treatment assignment) is randomly chosen, let alone that the distribution it was drawn from is known to the researcher (a point discussed in Manski and Pepper 2018, pages 234–235). Consequently, the prior conventional definition of identification in finite populations does not apply to such cases.
To address these two problems, we develop a different, formal definition of identification for a parameter in a finite population setting. Our definition is analogous to the usual definition of identification in infinite population settings (e.g., Manski 2003, Lewbel 2019, Molinari 2020): The finite population identified set for a parameter (such as the average treatment effect, ATE) is the set of parameter values consistent with (a) the finite matrix of data the researcher actually has and (b) any assumptions that they make.
Our definition directly addresses the two problems with the prior conventional definition described above: First, our definition does not require any knowledge about how treatment is assigned. Hence it can be applied in a wider variety of settings. Second, it uses actually observed data, rather than hypothetical data that is impossible to observe. Relatedly, it does not require any kind of asymptotics or hypothetical infinite populations. Consequently, any parameter which cannot be written as a known function of the actually observed data will typically be partially identified. The size of the identified set, however, will generally vary with the assumptions and data available, as in the usual infinite population definition of identified sets.
As a leading example of our identification analysis, we show that the average treatment effect in a randomized experiment is generally partially identified in finite populations. [^5] For a recent survey of infinite population approaches to randomized experiments, see Bai et al. (2025b). This conclusion arises from a subtle distinction between the finite and infinite population settings: Randomization guarantees that the treatment and control groups are exactly balanced in an infinite population, but not in finite populations. Since randomization cannot guarantee any level of balance between the treatment and control groups in a finite population, it does not have any identifying power.
Nonetheless, we show how to use our finite population identified sets combined with assumptions on treatment assignment to obtain a new way to quantify uncertainty in finite populations, which we call design-based sensitivity analysis. This is our second main contribution.
Overall, our method yields uncertainty intervals that can be interpreted as (1) identified sets, (2) robust Bayesian credible sets, or (3) uniform frequentist design-based confidence sets. This approach is non-asymptotic, which allows it to be reliably applied in small populations. And it allows researchers to flexibly impose additional assumptions on unobserved values, transparently quantifying the impact of these assumptions on conclusions. We summarize our approach in section 1.1.
In particular, focusing on their frequentist interpretation, our uncertainty intervals avoid several limitations of the confidence intervals that are currently available in the literature: (i) They do not rely on large-$N$ asymptotics, which can perform poorly if $N$ is small (e.g., $N=10$) and (ii) they allow for heterogeneous treatment effects. They are also valid uniformly over the set of all possible matrices of potential outcomes consistent with the assumptions (an important feature in partially identified settings; see Canay and Shaikh 2017). Furthermore, unlike the confidence intervals in the literature, our uncertainty intervals have both Bayesian and identification-based interpretations. This allows empirical researchers to choose their most preferred interpretation. This interpretational flexibility is one benefit obtained from a quantification of uncertainty that begins with identification. In particular, our uncertainty intervals are not based on hypothesis test-inversion, but are instead based on ideas from partial identification.
@@
As mentioned above, we develop this approach in the context of randomized experiments with a binary treatment, where the goal is to quantify the uncertainty in the average treatment effect due to missing potential outcomes. We extend the analysis in three main directions: In section 5 we give a new motivation for examining covariate balance. Specifically, by making an assumption that explicitly links covariates to potential outcomes, we show that observed imbalances in covariates have identifying power for the unobserved realized imbalance in potential outcomes, which allows researchers to obtain tighter bounds on ATE. In section 6 we study instrumental variable based analysis under one-sided noncompliance. And in section 7.1 we discuss uncertainty due to non-sampled units.
@@ 1.1 Our Approach to Quantifying Uncertainty in Experiments: A Practitioner’s Guide
Here we briefly sketch our approach, with an eye toward practical take-aways. We focus on causal inference here; see appendix LABEL:sec:sampleSurveyGuide for a parallel exposition in the context of survey sampling. Consider the population of six units in table 1 (left). Here $Y_{i}(1)$ and $Y_{i}(0)$ denote potential outcomes and $X_{i}\in\{0,1\}$ denotes realized treatment. The true ATE is 0.55, the difference between the average of $Y_{i}(1)$ over $i=1,\ldots,6$ and the average of $Y_{i}(0)$ over $i=1,\ldots,6$. The data, shown in the right table, only reveals half of the potential outcomes, however. Suppose we know that all outcomes must be between 0 and 1. Then without further assumptions, all we can say for sure about ATE is that it is in the set
@@
Finally, in step 3, we recommend combining these two plots as follows: For any desired ex ante probability $1-\alpha\in(0,1)$, use the bottom plot to find the smallest magnitude of imbalance that occurs with at least $100(1-\alpha)$% ex ante probability; denote this value by $K(\alpha)$. For example, if we set $1-\alpha=0.9$ then the figure shows us that $K=18.9$ satisfies $\underline{p}(K)=1-\alpha$ (see the dotted lines). We can then move to the top plot to read off bounds on ATE for this value of the sensitivity parameter, as shown in the dotted lines. In this case the bounds are about [6,22][-6,22]. Figure 1(b) shows these bounds for all values of $1-\alpha$. [^6] Note: For $1-\alpha$ close to zero but strictly positive, the bounds in figure 1(b) are non-singleton. This arises since exact balance is impossible in the worst case completions of table 1; this is why the plot of $\underline{p}$ in figure 1(a) is flat at zero up to a strictly positive value of $K$. We discuss this point further in section 2.3. We show that these uncertainty intervals have three interpretations:
@@
    Third, for any fixed $1-\alpha$, we show that these bounds are design-based confidence intervals. That is, across repeated re-assignments of treatment, these bounds will contain the true ATE at least $100(1-\alpha)$% of the time.
@@ 1.2 Related Literature
Our paper relates to a variety of literatures, but for brevity we only discuss the most closely related work, rather than attempt a comprehensive survey. Manski and Pepper (2018) derive identified sets in an explicitly finite population setting. Like them, we also consider the identifying power of bounded variation type assumptions. A main difference is that they focus on a setting with observational data whereas we primarily consider randomized experiments. Our focus on experiments allows us to use randomization itself to calibrate the magnitude of the bounded-variation sensitivity parameter $K$. Like us, Greenland and Robins (1986) also describe the problem of drawing conclusions in finite populations as an identification problem which must be solved by making assumptions about unobserved values of variables. They formally show how exact balance assumptions yield point identification. They then informally explain that “if we randomize…when both samples are large, random differences will in probability be small” (page 415). Many other papers make similar informal remarks, like Lindley (1980, page 590) and Royall and Pfeffermann (1982, page 480). Relatedly, the distinction between ex ante and ex post balance is well known; Greenland (1990) gives a particularly clear discussion, where he considers a randomized experiment with $N=2$. Greenland and Robins (2009) more recently discuss this distinction, and conclude that “randomization (or more generally, ignorability) does not impose “no [ex post] confounding”…rather, it provides…a randomization-based (“objective”) derivation of a prior…that applies after allocation as well as before, and becomes more narrowly centered around zero as the sample size increases. This is a key post-allocation benefit of randomization” (page 5, emphasis in the original). That paper as well as the rest of this prior literature, however, is largely verbal discussion. One of our main contributions is to build on those observations to provide a new, detailed, and formal approach to identification and quantifying uncertainty in finite populations.
@@
Finally, many recent papers on causal inference study explicitly finite populations, including Li and Ding (2017), Ding et al. (2017), Aronow and Samii (2017), Ding (2017), Athey et al. (2018), Kang et al. (2018), Abadie et al. (2020, 2023), Hong et al. (2020), Wu and Ding (2021), Eckles et al. (2020), Imbens and Menzel (2021), Zhao and Ding (2021), Sävje (2021), Bojinov et al. (2021), Xu (2021), Xu and Wooldridge (2022), Athey and Imbens (2022), Roth and Sant’Anna (2023), Pollmann (2023), Wooldridge (2023), Startz and Steigerwald (2023), Sancibrián (2024), Borusyak et al. (2025), Rambachan and Roth (2025), and Bai et al. (2025a), among many others. Also see Imbens and Rubin (2015) and Ding (2024) for book level surveys. This literature largely follows a template laid out by Neyman (1923) that jumps straight to deriving various frequentist properties, without doing any identification analysis. Two exceptions are Borusyak et al. (2025) and Rambachan and Roth (2025) which both discuss identification, based on the conventional definition, and do asymptotics-based inference (section 4.3 of Borusyak et al. 2025 also discuss exact inference under a constant effects assumption). Our finding that all potential outcomes are point identified under the conventional definition builds on ideas in Imbens (2018) and Chen et al. (2026), as we discuss in section 2.4. Rather than further studying identification, Chen et al. (2026) derive various frequentist testing and Bayesian updating implications of the conventional definition of identification, concluding that this conventional definition “translates to little, if any, practical learning” (page 14).
@@ 2 Finite Population Identification
We begin by describing our approach to identification: We define the finite population identified set in section 2.1. We then use this definition to derive identified sets for finite population average treatment effects under various assumptions in section 2.2, and we discuss the role of randomization for identification in section 2.3. We explain our concerns with the prior definition of identification in section 2.4.
@@ 2.1 The Finite Population Identified Set
Consider the standard potential outcomes model with a binary treatment. There are four components required to conduct an identification analysis:
1. The population: Suppose there are $N$ units in our population. Let $\mathcal{I}=\{1,\ldots,N\}$ be the set of indices for these units. Each unit $i\in\mathcal{I}$ is associated with the vector of numbers $(Y_{i}(1),Y_{i}(0),X_{i},W_{i})$, where $Y_{i}(1)$ and $Y_{i}(0)$ are potential outcomes, $X_{i}\in\{0,1\}$ is a realized binary treatment, and $W_{i}$ is a $d_{W}$-vector of covariates. The population is the $N\times(3+d_{W})$ matrix of all of these numbers. Let $\mathbf{P}^{\text{true}}\coloneqq(\mathbf{Y}(1),\mathbf{Y}(0),\mathbf{X},\mathbf{W})$ denote this population matrix. We use $\mathbf{P}$ to denote alternative possible values of this population matrix. Table 1 shows an example population.
2. Assumptions: We do not observe all elements of $\mathbf{P}^{\text{true}}$ because of the fundamental problem of causal inference. We typically make assumptions about its missing elements, the missing potential outcomes. Formalize these assumptions as the restriction that $\mathbf{P}^{\text{true}}\in\mathcal{P}$ where $\mathcal{P}$ is a known set of $N\times(3+d_{W})$ matrices. We give examples of $\mathcal{P}$ in section 2.2.
@@
4. The data: Finally, we must describe the data that is observed to the econometrician. To allow for sampling, let $\mathbf{S}=(S_{1},\ldots,S_{N})$ be a vector of indicators where $S_{i}=1$ if unit $i$ appears in the dataset, and zero otherwise. Let $\mathbf{Y}=(Y_{1},\ldots,Y_{N})$ where $Y_{i}\coloneqq Y_{i}(1)X_{i}+Y_{i}(0)(1-X_{i})$ for all $i\in\mathcal{I}$. Then $\mathbf{P}^{\text{data}}$ is the matrix of all rows of $(\mathbf{Y},\mathbf{X},\mathbf{W})$ with $S_{i}=1$. Formalize this construction via the function $\text{MakeData}(\mathbf{P},\mathbf{S})$, so that $\mathbf{P}^{\text{data}}=\text{MakeData}(\mathbf{P}^{\text{true}},\mathbf{S})$. [^7] We define the population at the time period after which treatment has been assigned but before units are sampled. This is standard in the identification literature (e.g., Manski 2003, ch. 7). One could modify the notation to further distinguish population values of potential outcomes and covariates from realized treatment assignment, as none of our concepts or procedures depend on this particular notational choice, but at the cost of additional notation throughout, especially since we mainly focus on the case where all units are sampled.
@@
This definition follows the usual description of the identified set as “the set of parameters…that are consistent with the model and the data” (Tamer 2010, page 184), but with close attention to what is meant by “data”: In a finite population, the “data” is defined by a finite dimensional matrix rather than a joint distribution of random variables (as in def. 3.1 on page 5324 of Matzkin 2007, for example; also see our appendix LABEL:sec:popDistributions for a related discussion). Our definition also allows for data on some units to be missing (i.e., not sampled, so that $S_{i}=0$); we discuss this further in section 7.1.
$\Theta_{I}$ has two key features: (1) It can always be computed with the data the analyst actually has at hand, and (2) like the usual definition of the identified set in an infinite population, it is guaranteed to contain the true parameter of interest, so long as the model is not false. We will revisit these properties below.